Circular No. 9121

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 D2 (WISE) A. Mainzer, Jet Propulsion Laboratory, reports that an object dis-

covered on 12-μm WISE satellite images (discovery observation tabulated below) shows a ‘hazy’ coma 20′′ in diameter with a 25′′ tail, with 24-μm images showing a coma diameter of ∼ 30′′. After posting on the ‘NEOCP’ webpage, P. Birtwhistle (Great Shefford, U.K., 0.40-m reflector) notes that his CCD images from Mar. 4.9 show a diffuse object of diameter 9′′ (total mag 19.5), possibly elongated in p.a. 75^{o}.

2010 UT α_{2000} δ_{2000} Observer
Feb. 25.68995 3^{h}22^{m}22.^{s}09 +71^{o}54′30′′.9 WISE

The available astrometry, the following parabolic orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2010-E12.

T = 2010 July 2.834 TT  ω = 143.^{o}791
Ω = 317.646 }2000.0
q = 3.66164 AU  i = 57.971

COMET 67P/CHURYUMOV-GERASIMENKO M. R. Combi, University of Michigan; J.-L. Bertaux and E. Quémerais,

Service d’Aéronomie, CNRS/UVSQ; and J. T. T. Mäkinen, Finnish Meteorological Institute, report that the Solar Wind Anisotropies (SWAN) camera on the Solar and Heliospheric Observatory (SOHO) spacecraft, located at the earth-sun L_{1} Lagrange point, made comet-dedicated observations of comet 67P in 2009 March and April. SWAN makes daily full-sky images of hydrogen Lyman-α. The following water-production rates were determined from the hydrogen Lyman-α brightness and distribution on each date, along with the heliocentric distance (r) of the comet: Mar. 2.53 UT, r = 1.248 AU, Q(H_{2}O) = 4.8 × 10^{27} molecules/s; 3.95, 1.248, 4.8; 5.36, 1.249, 4.9; 6.78, 1.250, 5.2; 8.20, 1.251, 4.3; 9.61, 1.253, 5.3; 11.03, 1.254, 5.8; 12.45, 1.256, 6.5; 13.86, 1.258, 6.0; 15.28, 1.261, 8.0; 16.70, 1.263, 8.9; 18.11, 1.266, 8.1; 19.53, 1.270, 7.6; 20.95, 1.273, 7.5; 22.35, 1.276, 6.8; 23.77, 1.280, 7.0; 25.19, 1.284, 5.3; 26.60, 1.288, 7.8; 28.02, 1.293, 8.4; Apr. 8.35, 1.335, 7.9; 9.77, 1.341, 11.1; 11.18, 1.348, 6.1; 12.60, 1.354, 4.3; 14.01, 1.361, 5.9; 15.42, 1.368, 5.3; 16.84, 1.375, 4.8; 18.26, 1.382, 3.3; 19.68, 1.389, < 2.5. The water production rates have uncertainties resulting from the calibration, background subtraction, and model parameters, as well as from the instrument noise and the formal model inversion. Taken together, these yield total uncertainties at a level of ∼ 30 percent.

2010 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green